OpenStudy (anonymous):

Jenna and Don collect autographs of baseball players. Jenna has five more autographs than don. If they have 19 autographs altogether, how many does each friend have?

OpenStudy (alfie):

x + (x+5) = 19 2x = 14 x= 7 Don = 7 Jenna = 7+5 = 12 12+7 = 19